# Evaluating and Improving Automatic Speech Recognition Systems for Korean Meteorological Experts

### v1

@@ Abstract
Evaluating and Improving Automatic Speech Recognition Systems
@@ 1 Introduction
Meteorological experts utilize diverse data sources for weather forecasting. The database systems they use are vast and complex, making it challenging for those without extensive system knowledge to access the necessary information. Natural language query systems can assist the experts by allowing them to input queries in natural language. The system translates these queries into database queries (e.g., SQL) to search the database (Zhong et al., 2017; Kim et al., 2020; Deng et al., 2022). This process enables forecasters to efficiently access and interpret complex databases without needing to write complex and structured queries themselves. Notably, Jo et al. (2023) developed an integrated search system for Korean weather data, enabling users to query extensive meteorological information using natural language. This system simplifies data retrieval and thereby improves operational efficiency.
@@ 3 Dataset Construction
This study aims to develop a specialized dataset for evaluating Automatic Speech Recognition (ASR) systems within the Korean meteorological domain. The dataset comprises 2,000 spoken queries derived from natural language questions. The construction process includes several key steps: utilizing the natural language question dataset from Jo et al. (2023), obtaining pronunciation guidelines for specialized terms from domain experts, collecting speech data from native speakers, and verifying the correctness of the recorded audio. The detailed data construction pipeline (section 3.1) and the analysis of the resulting dataset (section 3.2) are as follows.
@@ 3.1 Data Construction Pipeline
We began by utilizing the natural language question dataset released by Jo et al. (2023). This dataset includes a diverse range of questions relevant to the Korean meteorological domain, providing a robust foundation for our spoken query dataset. The dataset consists of two subsets: URL and SQL. We selected 1,300 questions from the URL subset and 700 questions from the SQL subset to ensure comprehensive coverage of various meteorological topics that forecasters typically seek. Note that the original dataset in Jo et al. (2023) contains natural language queries mapped to their corresponding structured queries (e.g., SQL). For our purposes, we used only natural language queries, focusing on how these questions would be articulated by users in real-world scenarios. For more details about natural language questions, please refer to the original paper of the dataset (Jo et al., 2023).
@@
Uttr. time (sec) 7.212.55 - min time 2.12 - max time 19.18 Avg. words 8.382.97 Avg. chars 32.0012.50 Unique words 2759 Absent ratio (%) 34.43
@@
Five native Korean speakers were recruited to record the spoken queries. Four were male, and one was female, all in their twenties. Each participant was given the pronunciation guidelines and detailed instructions on how to record the questions. Three participants recorded 500 queries each, while the other two recorded 200 and 300 queries, respectively. Participants were compensated at a rate of 16,000 KRW (approximately 11.6 USD) per 100 queries. The recordings were conducted in controlled environments to minimize background noise and ensure high audio quality. This phase resulted in a substantial collection of audio recordings that correspond to the natural language questions.
@@ 3.2 Dataset Analysis
Table 1 presents the overall statistics of our dataset, which consists of 2,000 spoken queries. The average utterance time is 7.21 seconds, ranging from 2.12 to 19.18 seconds, with each utterance containing an average of 8.38 words and 32.00 characters. The proportion of unique words in our dataset that are not found in the general domain ASR dataset (Bang et al., 2020) is 34.43%, highlighting the specialized and domain-specific nature of our dataset.
Table 2 shows the top 10 most frequently occurring words in both general and weather domain datasets. The general domain contains common conversational words, whereas the weather domain includes specialized terminology, underscoring the necessity of domain-specific datasets for improving ASR performance in meteorological applications.
General thing (거)    but (근데)    not (안)
that (그)    just (그냥)    I (내가)
such (그런)    uh (어)    what (뭐)
Weather surface (지상)   Standard Normals (평년값)
ocean (해양)   before (이전)   after (이후)
radar (레이더)   weather chart (일기도)
inform (알려줘)   satellite (위성)
Table 2: Ten most frequently occurred words in Korean ASR datasets in open-domain dialog (General) and weather (Weather) domains. The KsponSpeech (Bang et al., 2020) is used as a general domain dataset. Each English word is manually translated by the authors from the original Korean word presented in the bracket.
@@ 4.1 Evaluation Datasets
We use our Korean ASR dataset for the weather domain to evaluate different ASR models. The dataset comprises 2,000 spoken queries, which we randomly split into 500 samples for the development set and 1,500 samples for the test set. Additionally, we utilize the eval-clean test set from the KsponSpeech dataset (Bang et al., 2020) to evaluate the models in a general conversation domain. This dataset consists of 3,000 audio and their corresponding answer transcriptions. This approach allows us to compare models across both domain-specific and general ASR tasks.
@@ 4.3 Models and Training Datasets
Models
For our experiments, we utilized the multilingual Whisper model family (Radford et al., 2023), specifically leveraging four different model sizes: tiny (39M), small (244M), medium (769M), and large-v2 (1550M). These pre-trained models were evaluated in a zero-shot manner to assess their initial performance without additional training for our dataset. The target language and task in the prefix tokens of Whisper are set to Korean and transcribe, respectively.
@@
Model Params. KsponSpeechEval (3k) WeatherDev (0.5k) WeatherTest (1.5k)
CER WER sWER CER WER sWER CER WER sWER
Tiny 39M 32.94 55.26 51.59 35.12 78.74 58.14 34.06 77.61 56.95
Small 244M 16.88 33.56 27.51 16.92 47.45 29.39 17.00 48.56 29.61
Medium 769M 14.83 31.12 23.08 13.21 40.77 20.82 13.37 41.72 21.10
Large-v2 1550M 15.56 31.13 22.78 11.83 37.00 19.20 11.99 38.38 18.41
Table 3: Zero-shot ASR evaluation results of the Whisper model family. We used CER, WER, and sWER as evaluation metrics, where lower scores indicate better performance. The Weather dataset denotes our Korean weather-domain evaluation dataset. Whisper models of different sizes were evaluated. The lowest and the second-lowest scores are highlighted in bold and underlined, respectively.
Figure 3: Zero-shot qualitative results of pre-trained Whisper models with different model sizes. The answer queries in English are manually translated from the original Korean answer. Incorrect model predictions are highlighted with strikethrough. Omitted correct answers are highlighted with underline. Spacing errors are not highlighted for better visualization. Best viewed in color.
@@ 4.4 Implementation Details
Figure 4: Qualitative results of ASR models with different fine-tuning datasets. The answer queries in English are manually translated from the original Korean answer. ZS denotes the pre-trained Whisper large-v2 model in zero-shot evaluation. FT (General) and FT (Both) denote Whisper large-v2 models fine-tuned on KsponSpeech and both KsponSpeech and TTS-based Weather datasets, respectively. The indicators are the same as Fig. 3.
5 Results
Main Results
Table 3 shows the zero-shot evaluation results of Whisper models with different sizes in general and weather domains. We observed that model performance improved with increasing model size across both datasets. For example, in the KsponSpeech dataset, the Whisper large-v2 model achieved a CER of 15.56, compared to 32.94 for the Tiny model. Table 4 illustrates the impacts of fine-tuning datasets. Fine-tuning the Whisper large-v2 model on the KsponSpeech dataset significantly improved performance in the open-domain dialogue domain, reducing the CER to 10.48. However, this led to a decrease in performance in the weather domain, with a CER of 16.83. In contrast, fine-tuning on the TTS-based Weather dataset enhanced performance in the weather domain, achieving a CER of 6.56, while slightly reducing performance in the general domain. Combining fine-tuning on both KsponSpeech and the TTS-based Weather dataset provided the best overall results, showing substantial improvements over zero-shot performance in both domains, with a CER of 8.28 in the weather domain and 10.54 in the general domain. These findings suggest that increasing model size and using a combination of general and domain-specific training data effectively enhance ASR performance, offering meaningful gains in both general and specialized applications.
Fine-tune Data KsponSpeechEval WeatherTest General Weather CER WER CER WER 15.56 31.13 11.99 38.38 10.48 24.05 16.83 51.50 15.36 32.26 6.56 18.69 10.54 23.94 8.28 21.78
Table 4: Impacts of fine-tuning datasets. The Whisper-large-v2 model is used for the experiments. Models fine-tuned on Ksponspeech and TTS-based Weather datasets are marked with ✓in General and Weather, respectively. The indicators are the same as Table 3.
Qualitative Results
The qualitative results presented in Fig. 3 and Fig. 4 highlight the challenges faced by ASR models in accurately recognizing domain-specific terms. Fig. 3 shows results from pre-trained Whisper models of different sizes, while Fig. 4 includes models fine-tuned on the general domain (KsponSpeech) and both general and weather domains (KsponSpeech + TTS-based weather dataset). Zero-shot models often struggled with domain-specific terminology, frequently predicting incorrect but similar-sounding words. For instance, in the third example of Fig 3, the Tiny model mispredicted "입자계수기" ("ip-ja-gye-su-gi", "particle counter") as "입장해두기" ("ip-jang-hae-du-gi", admission). In the first example, the Large model incorrectly split the place name "백령도" ("Baeng-nyeong-do", "Baengnyeong island") into "100" ("baek", "one hundred") and "0도" ("0 do", "0 degrees"). In Figure 4, the fine-tuned model using both general and weather datasets showed improved accuracy in recognizing specialized terms like "낙뢰" (nak-roe, "lightning") and "하순" ("ha-sun", "late order"), demonstrating the effectiveness of domain-specific data augmentation in enhancing ASR performance.
@@ 6 Conclusion
In this study, we developed a specialized evaluation dataset to assess ASR models for Korean meteorological forecasters’ natural language queries. Our experiments demonstrated that pre-trained models often struggle to recognize domain-specific meteorological terms, highlighting the need for efforts to build reliable ASR models in specialized applications. The use of synthetic training samples for the weather domain showed some effectiveness in improving the accuracy of these specialized terms. These findings emphasize the importance of tailored evaluation datasets and domain-specific adaptations in developing robust ASR systems to support meteorologists in their forecasting tasks.

### v2

# Evaluating Automatic Speech Recognition Systems for Korean Meteorological Experts

@@ Abstract
Evaluating Automatic Speech Recognition Systems
@@ 1 Introduction
Meteorological experts utilize diverse data sources for weather forecasting. The database systems they use are vast and complex, making it challenging for those without extensive system knowledge to access the necessary information. Natural language query systems can assist the experts by allowing them to input queries in natural language. The system translates these queries into database queries (e.g., SQL) to search the database (Zhong et al., 2017; Kim et al., 2020; Deng et al., 2022). This process enables forecasters to efficiently access and interpret complex databases without writing complex and structured queries themselves. Notably, Jo et al. (2023) developed an integrated search system for Korean weather data, enabling users to query extensive meteorological information using natural language. This system simplifies data retrieval and thereby improves operational efficiency.
@@ 3 Dataset Construction
This study aims to develop a specialized dataset for evaluating Automatic Speech Recognition (ASR) systems within the Korean meteorological domain. The dataset comprises 5,500 spoken queries derived from natural language questions. The construction process includes several key steps: utilizing the natural language question dataset from Jo et al. (2023), obtaining pronunciation guidelines for specialized terms from domain experts, collecting speech data from native speakers, and verifying the correctness of the recorded audio. The detailed data construction pipeline (section 3.1) and the analysis of the resulting dataset (section 3.2) are as follows.
@@ 3.1 Data Construction Pipeline
We began by utilizing the natural language question dataset released by Jo et al. (2023). This dataset includes a diverse range of questions relevant to the Korean meteorological domain, providing a robust foundation for our spoken query dataset. The dataset consists of two subsets: URL and SQL. We selected 3,575 questions from the URL subset and 1,925 questions from the SQL subset to ensure comprehensive coverage of various meteorological topics that forecasters typically seek. Note that the original dataset in Jo et al. (2023) contains natural language queries mapped to their corresponding structured queries (e.g., SQL). For our purposes, we used only natural language queries, focusing on how these questions would be articulated by users in real-world scenarios. For more details about natural language questions, please refer to the original paper of the dataset (Jo et al., 2023).
@@
Num. samples 5,500
Uttr. time (sec) 7.052.55
- min time 0.92
- max time 29.98
Avg. words 7.594.34
Avg. chars 24.4915.62
Unique words 4,955
Absent ratio (%) 24.86
@@
Eleven native Korean speakers were recruited to record the spoken queries. Seven were male, and four were female, all in their twenties. Each participant was given the pronunciation guidelines and detailed instructions on how to record the questions. Participants were compensated at a rate of 16,000 KRW (approximately 11.6 USD) per 100 queries. The recordings were conducted in controlled environments to minimize background noise and ensure high audio quality. This phase resulted in a substantial collection of audio recordings that correspond to the natural language questions.
@@ 3.2 Dataset Analysis
Table 1 presents the overall statistics of our dataset, which consists of 5,500 spoken queries. The average utterance time is 7.05 seconds, ranging from 0.92 to 29.98 seconds, with each utterance containing an average of 7.59 words and 24.49 characters. The proportion of unique words in our dataset that are not found in the general domain ASR dataset (Bang et al., 2020) is 24.86%, highlighting the specialized and domain-specific nature of our dataset.
Table 2 shows the top 9 most frequently occurring words in both general and weather domain datasets. The general domain contains common conversational words, whereas the weather domain includes specialized terminology, underscoring the necessity of domain-specific datasets for improving ASR performance in meteorological applications.
General thing (거)    but (근데)    not (안) that (그)    just (그냥)    I (내가) such (그런)    uh (어)    what (뭐) Weather show (보여줘)   AWS (AWS) inform (알려줘)   standard normals (평년값) surface (지상)   highest temperature (최고 기온) ocean (해양)   radar (레이더)  weather chart (일기도)
Table 2: Nine most frequently occurred words in Korean ASR datasets in open-domain dialog (General) and weather (Weather) domains. The KsponSpeech (Bang et al., 2020) is used as a general domain dataset. Each English word is manually translated by the authors from the original Korean word presented in the bracket.
@@ 4.1 Evaluation Datasets
We use our Korean ASR dataset for the weather domain to evaluate different ASR models. The dataset comprises 5,500 spoken queries, which we randomly split into 500 samples for the development set and 5,000 samples for the test set. Additionally, we utilize the eval-clean test set from the KsponSpeech dataset (Bang et al., 2020) to evaluate the models in a general conversation domain. This dataset consists of 3,000 audio and their corresponding answer transcriptions. This approach allows us to compare models across both domain-specific and general ASR tasks.
@@ 4.3 Models and Training Datasets
Zero-shot Multilingual Models
For our experiments, we used the multilingual Whisper model family (Radford et al., 2023), evaluating four different model sizes: tiny (39M), small (244M), medium (769M), and large-v2 (1550M). These pre-trained models were assessed in a zero-shot manner to evaluate their performance without additional fine-tuning on our dataset. The target language and task in Whisper’s prefix tokens were set to Korean and transcribe, respectively. Additionally, for evaluation in the weather domain, we tested Google’s commercial speech-to-text API [^2] https://cloud.google.com/speech-to-text.
@@
Model Params. KsponSpeechEval (3k) WeatherDev (0.5k) WeatherTest (5.0k) CER WER sWER CER WER sWER CER WER sWER (1) Zero-shot Evaluation of Whisper model family (Radford et al., 2023) Tiny 39M 32.67 53.69 47.78 33.53 74.93 54.19 38.24 81.99 58.61 Small 244M 16.75 32.57 26.2 16.39 46.62 27.69 22.92 57.4 34.65 Medium 769M 14.33 29.73 21.21 12.65 38.67 19.31 19.26 50.38 26.54 Large-v2 1550M 14.66 29.97 20.75 11.78 36.31 19.03 18.35 46.74 25.46 (2) Whisper-Large-v2 finetuned on Different Datasets General 1550M 10.48 24.05 16.49 16.92 49.59 21.75 24.57 63.28 28.22 Weather 1550M 15.81 32.48 25.28 6.93 17.99 9.13 10.64 31.81 13.61 General+Weather 1550M 10.54 23.94 16.52 8.68 22.58 12.50 17.46 43.41 20.25 (3) Commercial API Google STT - - - - 20.38 57.09 29.45 23.45 71.89 31.10
Table 3: ASR evaluation results. We report CER, WER, and sWER, where lower values indicate better performance. The table consists of three model groups: (1) zero-shot Whisper models of varying sizes, (2) Whisper-large-v2 fine-tuned on different datasets, and (3) a commercial ASR system (i.e., Google STT). In Group (2), General refers to fine-tuning on KsponSpeech, and Weather denotes fine-tuning on the TTS-generated weather dataset. The lowest and second-lowest scores of each column are highlighted in bold and underlined, respectively.
@@
Post-processing with Unimodal LLMs
To evaluate the effectiveness of large language models (LLMs) in refining ASR transcriptions without utilizing acoustic information (Chen et al., 2023; Hu et al., 2024a), we conducted an experiment where GPT-4o-mini (Achiam et al., 2023) post-processed the top-1 transcription from Whisper-large-v2. Specifically, we provided the LLM with the single best output from the ASR model. The model was tested in both a zero-shot setting and a 20-shot setup using examples from the development set. We adopted the text prompt from (Chen et al., 2023) and tailored it to the Korean meteorological domain to ensure corrections aligned with domain-specific terminology. (i.e., "Below is the best-hypotheses transcribed from speech recognition system for Korean Meteological experts. Please try to revise it and write the response for the true transcription.") This experiment allows us to assess how effectively an LLM can improve ASR accuracy in a specialized domain when limited to a single-hypothesis input.
@@ 4.4 Implementation Details
Figure 3: Effect of LLM-based post-processing on ASR. We compare ASR performance under three conditions: without LLM-based grammatical error correction (w/o GEC), with zero-shot post-processing (0 Shot), and with a 20-shot setup (20 Shot).
5 Results and Analyses
We present the key findings from our ASR evaluation on the Korean meteorological dataset. Table 3 summarizes the results across different setups.
Zero-shot ASR models usually struggle with domain-specific terminology
Whisper models exhibit high CER and WER when transcribing meteorological queries, particularly for specialized terms. While larger models (e.g., Whisper-large-v2) reduce errors compared to smaller variants, they still struggle with domain-specific terminology, reinforcing the need for adaptation.
Fine-tuning improves general performance, but domain adaptation is crucial
Fine-tuning Whisper-large-v2 on KsponSpeech enhances general transcription accuracy, lowering CER and WER across datasets. However, performance gains on meteorological queries remain limited, suggesting that general-domain adaptation alone is insufficient for specialized terminology.
Synthetic data (TTS) effectively enhances domain adaptation
Training with TTS-generated weather-domain data significantly improves transcription accuracy for meteorological terms. The model fine-tuned on both KspSpeech and synthetic weather data achieves the lowest CER (8.68%) and WER (22.58%) on the test set, demonstrating the effectiveness of controlled synthetic augmentation.
LLM-based post-processing does not consistently improve ASR output
As shown in Figure 3, the LLM-based post-processing does not consistently improve meteorological term recognition. The CER, WER, and sWER remain higher than those of the zero-shot Whisper model, indicating that text-based refinement alone is insufficient for domain-specific transcription without additional acoustic training. While training a specialized GEC model on the target dataset or leveraging large-N hypothesis generation could potentially enhance performance (Chen et al., 2023; Hu et al., 2024b), these approaches would require a large-scale training corpus and significantly increase computational costs during inference.
Commercial ASR systems underperform in specialized tasks
Google’s STT model, a widely used commercial ASR system, yields the highest error rates (e.g., 20.38% on Weather-test set), performing worse than zero-shot Whisper-large-v2. This highlights the limitations of generic ASR systems in handling meteorological speech and underscores the necessity of domain-specific adaptation.
These results emphasize the importance of fine-tuning ASR models with domain-specific data. While general LLMs can aid transcription refinement, substantial performance gains require dedicated acoustic adaptation.
Answer 8월 30일 증기압과 기온 평년값은 뭐니 ZS 8월 30일 증기압과 기운 평량값은 뭐니 FTGeneral 8월 30일 기압과 기온 평년값은 뭐니 FTBoth 8월 30일 증기압과 기온 평년값은 뭐니 Google STT 8월 30일 증기압과 기온 평년 값은 뭐니 LLM20shot 8월 30일 증기압과 기온 평값은 무엇인가요 Answer 관측횟수 CALM 0 5m/s 순별 하순 211 인제 ZS 관측해수컴 0 5mf 순별 하순 211 인 FTGeneral 관측 해수 칸 영 5m 퍼셉커 순별 하순 211 인제11 인제 FTBoth 관측해서 CALM 0 5m순별 하순 2 1 1 인제 Google STT 관측에서 컴 0 5 m 버서커 순별 순 2 인제 LLM20shot 관측 0$\sim$5 m/s 순별 하순 211 인
Table 4: Qualitative Results Incorrectly recognized characters are marked with strikethrough, while missing words are indicated with a wavy underlinewavy underline. Space errors are omitted for better readability.
Case Study
To illustrate common ASR challenges, we examine qualitative examples from Table 4. One issue arises with numerical and meteorological term recognition. In a wind forecast transcription, "지상 시계열 바람장미 풍속계급별 판촉횟수 CALM 0 5m/s 순별 하순 211 인제" ("Ground-level wind rose classified by wind speed, CALM 0–5m/s, late-stage 211 Inje"), zero-shot Whisper and Google STT misinterpret "CALM", a standard meteorological term for low wind speeds, distorting critical numerical values. Another case involves unintended style shifts. In "8월 30일 증가압과 기온 평년값은 뭐니" ("What are the historical averages for pressure and temperature on August 30?"), the post-processed output rephrases it as "8월 30일 증가압과 기온 평년값은 무엇인가요?", making the tone more formal. While this improves fluency, it may be undesirable for contexts requiring colloquial speech. These cases highlight two key challenges: (1) accurately preserving domain-specific terms and numerical values, and (2) balancing grammatical refinement with maintaining the intended speech style. Addressing these requires targeted domain adaptation and context-aware correction strategies.
@@ 6 Conclusion
This work constructs a specialized evaluation dataset for assessing ASR models in the Korean meteorological domain, addressing the lack of domain-specific benchmarks. Our results show that pre-trained ASR models struggle with meteorological terminology, and while fine-tuning on general-domain data improves accuracy, domain adaptation remains essential for handling specialized vocabulary. While our dataset provides a valuable benchmark, future work should expand it to cover more diverse speech conditions, such as spontaneous and noisy environments, to better reflect real-world forecasting scenarios. By making this dataset available, we aim to drive further advancements in domain-adapted ASR systems for meteorological applications.
@@ Appendix A Pronunciation Guideline Details for Speech Data Collection
Appendix B Qualitative Results
We present further prediction results of different open-source ASR models on Figures 5 and 6.
Figure 5: Qualitative results of pre-trained Whisper models with different model sizes. The answer queries in English are manually translated from the original Korean answer. Incorrect model predictions are highlighted with strikethrough.
Figure 6: Qualitative results of Whisper models finetuned on different datasets. All indicators are the same with Figure 5.